OpenStudy (anonymous):

Find the slope of the line 5x-2y=3 is the slope 5/2?

OpenStudy (anonymous):

5x-2y=3 -2y=-5x+3 Divide all by negative 2 y= 5/2 x -3/2 The slope = 5/2x